# Unstructured Evidence Attribution for Long Context Query Focused Summarization

### v1

@@ Abstract
Large language models (LLMs) are capable of generating coherent summaries from very long contexts given a user query. Extracting and properly citing evidence spans could help improve the transparency and reliability of these summaries. At the same time, LLMs suffer from positional biases in terms of which information they understand and attend to, which could affect evidence citation. Whereas previous work has focused on evidence citation with predefined levels of granularity (e.g. sentence, paragraph, document, etc.), we propose the task of long-context query focused summarization with unstructured evidence citation. We show how existing systems struggle to generate and properly cite unstructured evidence from their context, and that evidence tends to be “lost-in-the-middle”. To help mitigate this, we create the Summaries with Unstructured Evidence Text dataset (SUnsET), a synthetic dataset generated using a novel domain-agnostic pipeline which can be used as supervision to adapt LLMs to this task. We demonstrate across 5 LLMs of different sizes and 4 datasets with varying document types and lengths that LLMs adapted with SUnsET data generate more relevant and factually consistent evidence than their base models, extract evidence from more diverse locations in their context, and can generate more relevant and consistent summaries.
@@ 1 Introduction
At the frontier of the capabilities of natural language processing (NLP) systems such as large language models (LLMs) is the ability to handle long contexts such as books, sets of research papers, and long legal documents, and summarize them based on user queries Koh et al. (2023); Su et al. (2024); Beltagy et al. (2020); Reid et al. (2024). The difficulty of this task lies in the need to attend to relevant information in the source document(s) given a query and simultaneously derive coherent, factually consistent, and distilled insights Goldman et al. (2024). While LLMs have achieved much progress on this Edge et al. (2024), people prefer to use traditional retrieval sources (e.g., search engines) for critical queries due to the need for transparency and provenance Worledge et al. (2024). While progress has been made on structured evidence citation (i.e., with fixed levels of granularity such as sentences and documents) Li et al. (2023), in order to improve the flexibility and explainability of long-context query focused summaries we propose to study the task of unstructured evidence citation.
Figure 1: We study query focused summarization with unstructured evidence citation from long contexts
Within this, we explore two key barriers to achieving good performance with reliable summaries. First, LLMs have positional biases Liu et al. (2024c); Ravaut et al. (2024), focusing on the earlier and later tokens in their input context Zhang et al. (2024b). This can potentially bias which evidence a model selects to support a summary. Second, fine-tuning has benefits in terms of inference efficiency Wu et al. (2024), and in some cases can outperform inference interventions Zhang et al. (2024a); Liu et al. (2022), but would require a large dataset with specialized examples of long documents, queries, extracted evidence, and summaries which cite this evidence. Creating such a dataset requires an extensive amount of time, money, and expertise Asai et al. (2024); Laban et al. (2024); Santosh et al. (2024).
To address these issues, we provide the first study on unstructured (i.e., no fixed level of granularity) evidence citation in long-context query-focused summarization. While attribution has been investigated in recent works Laban et al. (2024); Asai et al. (2024); Li et al. (2023), we present a novel and more challenging scenario where a model must extract unstructured text spans from its context to use as supporting information for its summary. We show for base models that extracting and citing unstructured evidence is challenging, and that evidence is often lost-in-the-middle. To help alleviate this, we present a fine-tuning approach based on an inductively generated synthetic dataset called the Summaries with Unstructured Evidence Text dataset (SUnsET). Across 5 different models and 4 different datasets (single- and multi-document), we demonstrate that fine-tuning on SUnsET data can help mitigate lost-in-the-middle, increase citation accuracy and coverage, and improve summary quality. We release SUnsET and our generation code to the public for further study. [^1] https://github.com/dwright37/unstructured-evidence-sunset
2 Challenges in Query Focused Long Context Summarization
@@
Query focused, long context summarization requires a model to be able to simultaneously ingest a large number of context tokens (possibly from multiple documents), retrieve and attend to relevant information in this context given a query, and integrate this information into a factually consistent and relevant summary. LLMs, with their increasingly large context sizes, have proven to be particularly adept at performing this task Zhang et al. (2024a); Edge et al. (2024); Russak et al. (2024). Yet, a number of challenges remain, both in dealing with long contexts and with producing query-focused summaries Li et al. (2024); Russak et al. (2024); Bai et al. (2024); Liu et al. (2024c); Shaham et al. (2023); Ravaut et al. (2024); Laban et al. (2024); Worledge et al. (2024); Ji et al. (2023). The main focus of our work is on evidence attribution Laban et al. (2024); Worledge et al. (2024); Li et al. (2023) and its relation to the lost-in-the-middle problem Liu et al. (2024c); Ravaut et al. (2024).
@@ 2.1 Evidence Attribution
Though LLMs can generate convincing summaries, in practice people often prefer to acquire information through media where provenance is fully transparent (for example, a normal search engine) Worledge et al. (2024). Improving the ability of LLMs to generate both relevant summaries and provide accurate attributions has the potential to help improve their usefulness, transparency, and trustworthiness. Recent work has started to explore this direction, including SummHay Laban et al. (2024) and OpenScholar Asai et al. (2024). However, most works focus on structured attribution with fixed levels of granularity (e.g., sentences, paragraphs, or documents) Li et al. (2023). To the best of our knowledge, we provide a first study on unstructured evidence citation in long context query focused summarization, which is more flexible than the fixed-granularity approach.
@@ 2.2 Lost-in-the-Middle
LLMs suffer from positional preferences in their learned attention Liu et al. (2024c), oftentimes preferring early or late tokens in their context Zhang et al. (2024b). While this problem was originally demonstrated on retrieval-augmented-generation (RAG) tasks with explicit answers such as question answering, follow-up work has shown its persistence in more abstractive tasks such as summarization Ravaut et al. (2024) and query focused multi-document summarization Laban et al. (2024). A number of solutions have been proposed, most of which rely on manipulating either the positions of tokens in the context or the positional embeddings of LLMs in order to remove their intrinsic bias Wang et al. (2025); He et al. (2024); Zhang et al. (2024b). We further explore and document this problem at the level of unstructured evidence citation, demonstrating how evidence is extracted unevenly across documents, and how this problem can be mitigated using purely synthetic data.
3 Learning to Cite and Summarize
Our task is: given a query about a long input consisting of one or more documents, generate a response to the query which cites unstructured evidence from the input. We differentiate ourselves from previous work on summarization with attribution Laban et al. (2024); Asai et al. (2024); Li et al. (2023) by requiring a model to extract unstructured text as evidence with no predefined levels of granularity. This evidence must also be relevant and consistent with both the document and the summary sentences. While more challenging, this enables flexible and explainable summary generation, and citations can easily be generated in both multi-document and single-document settings.
@@ 3 Learning to Cite and Summarize
1]]]] Writing the unwritable requires a recognition of the limitations of language, and a willingness to push against those boundaries.
@@
To tackle this we create SUnsET, a synthetic dataset which is generalizable across domains, using a novel inductive generation pipeline. We then use this data to fine-tune models using adapters Houlsby et al. (2019) to improve unstructured evidence citation and summary quality, as well as mitigate the lost in the middle problem. For the latter, previous work has shown that fine-tuning with data augmentation (e.g., shuffling documents Zhang et al. 2024b) can help achieve this. Given this, we construct SUnsET so that documents are modular: documents are broken down into discrete sections, so that data augmentation through shuffling document sections (thus shuffling global structure) is possible. We first present the inductive pipeline approach used to generate SUnsET, followed by our two fine-tuning schemes which use low-rank adapters (LoRA) Hu et al. (2022) trained on different views of SUnsET.
SUnsET Base Base w/ Titles
Method Titles Questions Summaries Documents Titles Questions Summaries Documents Titles Questions Summaries Documents
Moving Avg. TTR 0.816 0.751 0.836 0.820 0.387 0.670 0.797 0.350 0.588 0.631 0.778 0.352
Avg. Cosinse Dist. 0.780 0.806 0.733 0.682 0.425 0.725 0.716 0.042 0.607 0.660 0.610 0.040
Avg. Length (in words) 5.44 13.45 226.5 3767.4 6.65 9.85 23.79 474.8 5.76 10.21 24.45 433.8
Table 1: Statistics and diversity metrics of synthetically generated data.
3.1 Generating SUnsET
Recent works have demonstrated the promise of generating large scale synthetic datasets for fine-tuning task specific models Ziegler et al. (2024); Honovich et al. (2023); Wang et al. (2023); Chen et al. (2024); Xu et al. (2024) with the help of strong LLMs such as GPT-4 OpenAI (2023). Inspired by this, we propose SUnsET, a generic, domain-agnostic dataset generated from a novel inductive synthetic data pipeline that allows us to fine-tune downstream models to generate relevant and consistent query focused summaries from long-contexts with unstructured evidence citations.
Our pipeline generates datasets composed of long documents paired with queries and long-form answers to those queries. Each summary includes inline citations that reference relevant unstructured text spans in the original document. We make several design decisions intended to overcome known problems in synthetic data generation, including the potential for low diversity Honovich et al. (2023); Wang et al. (2023) and labeling errors Chen et al. (2024). This includes taking a six stage pipeline approach which generates synthetic data inductively, and validation steps which refine summaries, refine evidence, and reject bad summaries and evidence.
The full generation process is described in Figure 2, with prompts provided in Appendix A. Diversity in document topic and type is accomplished by first generating diverse document titles, which seed the subsequent steps of generation. We inductively build up each document, starting with the queries, summaries, and evidence passages. When generating evidence, each evidence passage is assigned to a section in the document so that evidence can be distributed precisely. The summaries, queries, and assigned evidence are then used as context from which each section of the document is generated, one section at a time. This makes documents modular, which we take advantage of during training to study evidence positional biases. Following this, the queries, summaries, and evidence are refined in order to fully reflect the final document. Finally, we filter the summaries and evidence by prompting GPT 4o mini to predict if the summaries fully address the query and are fully supported by the document (see Figure 3 for an example).
To validate the pipeline, we additionally generate two baseline datasets. The first is generated by combining all the steps listed in Figure 2 into a single prompt. The second includes a control where we enforce that no repeated titles are generated (see Figure 17 in Appendix A). We compare each dataset using samples of 100 documents along dataset diversity metrics (average type-token ratio (TTR)Bestgen (2023), embedding cosine distance, and average word length) in Table 1. Baseline non-pipelined approaches produce shorter documents and shorter summaries, and these documents and summaries tend to be much more similar to each other than those generated using our pipeline.
@@ 3.2 Training Complementary Adapters
Previous work has demonstrated that altering the position embeddings of LLMs either directly or through fine-tuning can help to overcome positional biases Hsieh et al. (2024); Zhang et al. (2024b). One of the benefits of SUnsET documents is that they are highly modular, as the generated documents have both global coherence at the level of the full document and local coherence at the level of discrete sections. Given this, we experiment with position-aware and position-agnostic training in order to observe their impact on evidence selection and quality, as well as summary quality.
For position-aware training, we concatenate all the document sections together in their natural order to construct the context, while for position-agnostic training, we shuffle the document sections before concatenating them, thus randomizing the global structure of the position embeddings while maintaining the local structure. This gives us two adapters for each model in our experiments. The prompt we use for training is provided in Appendix A Figure 18, and all training is performed using supervised fine-tuning on SUnsET data using LoRA Hu et al. (2022).
@@ 3.3 Summarizing with Unstructured Evidence
Method Exact Match 50% Match # Evidence
Llama 3.2 1B  0.0 35.71     14
   + Standard   7.69 43.26   208
   + Shuffled   5.15 22.68     97
Llama 3.2 3B 25.57 90.11 1345
   + Standard 52.77 85.62 3720
   + Shuffled 32.99 74.07 2337
Llama 3.1 8B 43.93 83.12 3412
   + Standard 78.36 97.21 4690
   + Shuffled 54.53 88.51 4684
Mistral Nemo 2407   5.48 66.13   310
   + Standard 82.20 97.29 2107
   + Shuffled 72.38 95.76 1959
Mixtral 8x7B   5.79 91.25 3452
   + Standard 33.82 90.47 4208
   + Shuffled 29.29 90.74 4288
GPT-4o-mini 11.06 96.32 8159
Table 2: Hallucination rates for evidence extraction. We directly measure exact string match (i.e. when the evidence sentence exactly appears in the context) as well as 50% overlap between the extracted evidence and the longest common substring in the context.
SLTS LASS SMHM SQBM
Method RelF1 ConF1 RelF1 ConF1 RelF1 ConF1 RelF1 ConF1
Llama 3.2 1B 0.00 0.00 0.94 1.06 0.00 0.00 0.23 0.18
+ Standard 0.63 0.53 4.80 4.56 0.00 0.00 1.84 1.68
+ Shuffled 0.48 0.26 2.83 2.84 0.00 0.00 0.00 0.00
Llama 3.2 3B 11.21 10.16 15.08 14.64 8.64 8.75 12.37 12.99
+ Standard 36.19 25.12 43.98 40.64 37.73 39.03 37.16 34.39
+ Shuffled 23.38 15.33 36.19 31.26 32.73 33.46 31.36 26.73
Llama 3.1 8B 17.21 15.15 31.17 30.65 34.18 37.96 32.08 32.85
+ Standard 35.21 25.34 52.64 47.79 56.82 57.50 45.26 41.13
+ Shuffled 29.36 20.65 49.90 44.19 54.79 54.27 39.53 36.17
Mistral Nemo 2407 2.75 2.37 5.34 4.58 10.37 10.25 5.67 5.36
+ Standard 34.24 24.45 38.21 36.88 23.54 25.13 7.15 7.56
+ Shuffled 32.52 22.84 39.94 38.57 21.58 23.23 4.65 4.08
Mixtral 8x7B 24.45 19.15 39.48 40.08 44.01 43.44 25.97 25.61
+ Standard 30.54 25.11 38.27 38.08 48.71 51.85 38.37 38.59
+ Shuffled 32.87 25.86 44.13 44.48 46.67 49.09 39.65 41.89
GPT 4o Mini 42.62 36.23 59.48 53.96 64.99 60.14 37.65 33.11
Table 3: Relevance and consistency of evidence sentences with respect to their citances. Relevance and consistency are measured using an autorater (GPT-4o-mini) Liu et al. (2023) based on previously validated prompts Liu et al. (2024b). We follow a similar setup to Laban et al. (2024); Asai et al. (2024) where we measure citation precision and recall in order to calculate an overall F1 score for both relevance and consistency. S indicates single document tasks, M indicates multi-document. SQ is SQuALITY, LAS is LexAbSumm, SMH is SummHay, and SQB is ScholarQABench

### v2

@@ Abstract
Large language models (LLMs) are capable of generating coherent summaries from very long contexts given a user query, and extracting and citing evidence spans helps improve the trustworthiness of these summaries. Whereas previous work has focused on evidence citation with fixed levels of granularity (e.g. sentence, paragraph, document, etc.), we propose to extract unstructured (i.e., spans of any length) evidence in order to acquire more relevant and consistent evidence than in the fixed granularity case. We show how existing systems struggle to copy and properly cite unstructured evidence, which also tends to be “lost-in-the-middle”. To help models perform this task, we create the Summaries with Unstructured Evidence Text dataset (SUnsET), a synthetic dataset generated using a novel pipeline, which can be used as training supervision for unstructured evidence summarization. We demonstrate across 5 LLMs and 4 datasets spanning human written, synthetic, single, and multi-document settings that LLMs adapted with SUnsET generate more relevant and factually consistent evidence with their summaries, extract evidence from more diverse locations in their context, and can generate more relevant and consistent summaries than baselines with no fine-tuning and fixed granularity evidence. We release SUnsET and our generation code to the public. [^1] https://github.com/dwright37/unstructured-evidence-sunset
@@ 1 Introduction
Figure 1: Summarization with unstructured evidence requires a model to retrieve spans of any arbitrary length from the context to support individual sentences in the summary. Example given from Llama 3.1 8B trained on our dataset (SUnsET).
@@
At the frontier of the capabilities of natural language processing (NLP) systems such as large language models (LLMs) is the ability to handle long contexts such as books and research papers, and summarize them based on queries Koh et al. (2023); Su et al. (2024); Beltagy et al. (2020); Reid et al. (2024). While LLMs have progressed much on this Edge et al. (2024), people prefer traditional retrieval sources (e.g., search engines) for critical queries due to transparency and provenance Worledge et al. (2024). Citing evidence in the summary addresses this, with prior work first segmenting the context into spans at fixed levels or granularity (e.g., sentences or documents, see Li et al. 2023) and having models select evidence from among these segments to support the summary. As has been noted both in work on multi-document summarization Ernst et al. (2024); Xiao (2023) and automated fact checking Wan et al. (2021), this approach is suboptimal for acquiring the most salient text in the context to support the summary, resulting in either too much or not enough information. In order to improve the precision of evidence in long-context query focused summarization (LCQFS), we propose to study unstructured evidence citation, where any span of arbitrary length within the context can be used as evidence.
In the unstructured evidence setup, a model must first copy spans from the context and subsequently use those spans as evidence in the summary (see Figure 1). As we will show, simply prompting LLMs to perform this task with no other intervention leads to poor performance. Thus, we need to adapt models, e.g. through fine-tuning or in-context learning. For this, no suitable training data exist which consists of examples of long documents, queries, summaries, and extracted evidence pointing to arbitrary spans in the documents. Based on the size and cost of other datasets for LCQFS Asai et al. (2024); Laban et al. (2024); Santosh et al. (2024), this would take an extensive amount of time, money, and expertise to create manually.
To address this, we present a synthetic dataset called the Summaries with Unstructured Evidence Text dataset (SUnsET). SUnsET is generated using a novel pipeline, resulting in long documents paired with queries, summaries, and evidence spans. We show that the data in SUnsET are high quality and diverse, comparable to human written data. Using SUnsET, we perform experiments across 5 models and 4 test datasets (including single- and multi-document, human and synthetic data), leading to the following findings: 1) for base LLMs with no fine tuning, extracting and citing unstructured evidence is challenging, and evidence is often lost-in-the-middle; 2) training on documents with shuffled structure (facilitated by SUnsET) can help mitigate lost-in-the-middle, and 3) learning to cite unstructured evidence improves citation accuracy and coverage over fixed-granularity evidence, and additionally improves summary quality.
In sum, our contributions are:
  • A synthetic dataset (SUnsET) generated using a novel pipeline
  • The first study on unstructured evidence citation for LCQFS, demonstrating that models adapted with SUnsET produce higher quality evidence and summaries than baselines
  • An analysis of and method to reduce the lost-in-the-middle problem with unstructued evidence
2 Challenges in LCQFS
@@
LCQFS requires a model to be able to simultaneously ingest a large number of context tokens (possibly from multiple documents), retrieve and attend to relevant information in this context given a query, and integrate this information into a factually consistent and relevant summary. LLMs, with their increasingly large context sizes, have proven to be particularly adept at performing this task Zhang et al. (2024a); Edge et al. (2024); Russak et al. (2024). Yet, a number of challenges remain, both in dealing with long contexts and with producing query-focused summaries Li et al. (2024); Russak et al. (2024); Bai et al. (2024); Liu et al. (2024b); Shaham et al. (2023); Ravaut et al. (2024); Laban et al. (2024); Worledge et al. (2024); Ji et al. (2023); Ernst et al. (2024). The main foci of our work are evidence attribution Laban et al. (2024); Worledge et al. (2024); Li et al. (2023); Ernst et al. (2024); Fierro et al. (2024) and evidence being lost-in-the-middle Liu et al. (2024b); Ravaut et al. (2024), described next.
Fixed-Granular Single Sentence Citation: Summary Snippet: …[48] explains that the legend of the Ghost Ship is often told by space men as a cautionary tale…. Evidence: [48] He had heard it spoken of in whispers by drunken space men and professional tellers of fairy tales.   Unstructured Citation: Summary Snippet: …he, like the ship’s former crew, is doomed to wander in space, never able to return to Earth, a haunting reminder of what he has lost and what he can never have [2]… Evidence: [2] Doomed for all eternity to wander in the empty star-lanes, the Ghost Ship haunts the Solar System that gave it birth. And this is its tragedy, for it is the home of spacemen who can never go home again.
Figure 2: Examples of fixed-granular and unstructured evidence generated by models in our study. Fixed granular citations may include irrelevant or not enough information to support their citing sentences. Unstructured evidence allows for more flexible and precise evidence.
@@ 2.1 Evidence Attribution
Improving the ability of LLMs to both generate relevant summaries and provide accurate attributions has the potential to help improve their usefulness, transparency, and trustworthiness. Recent work has started to explore this direction for LCQFS, including SummHay Laban et al. (2024) and OpenScholar Asai et al. (2024). However, most works focus on fixed-granularity evidence (e.g., spans, sentences, paragraphs, or documents, Li et al. (2023)). Being able to flexibly cite evidence of any arbitrary length can lead to higher quality summaries which use precise pieces of evidence from the context Wan et al. (2021); Ernst et al. (2024); Xiao (2023), as opposed to full documents which contain irrelevant information or individual sentences which may contain not enough information (see e.g., Figure 2). To the best of our knowledge, we provide a first study on unstructured evidence citation in LCQFS with LLMs.
@@ 2.2 Lost-in-the-Middle
LLMs suffer from positional preferences in their learned attention Liu et al. (2024b), oftentimes preferring early or late tokens in their context Zhang et al. (2024b). While this problem was originally demonstrated on retrieval-augmented-generation (RAG) tasks with explicit answers such as question answering, follow-up work has shown its persistence in more abstractive tasks such as summarization Ravaut et al. (2024) and query focused multi-document summarization Laban et al. (2024). A number of solutions have been proposed, most of which rely on manipulating either the positions of tokens in the context or the positional embeddings of LLMs in order to remove their intrinsic bias Wang et al. (2025); He et al. (2024); Zhang et al. (2024b). We explore and document this problem at the level of unstructured evidence citation, demonstrating how evidence is extracted unevenly across documents, and how this problem can be mitigated using purely synthetic data.
3 Learning to Use Unstructured Evidence
Our task is: given a query about a long input consisting of one or more documents, generate a response to the query which cites arbitrary length text spans from the input. This introduces challenges over the fixed-granularity case Laban et al. (2024); Asai et al. (2024); Li et al. (2023), as targeted, precise evidence spans must be accurately copied from the context which are relevant and consistent with the summary sentences. While challenging, this can lead to summaries with more accurate and supportive evidence (Ernst et al. 2024).
@@ 3 Learning to Use Unstructured Evidence
Large scale synthetic datasets are useful for fine-tuning task specific models at a lower cost than manual annotation Ziegler et al. (2024); Honovich et al. (2023); Wang et al. (2023); Chen et al. (2024); Xu et al. (2024). To train LLMs to use unstructured evidence, we create SUnsET, a synthetic dataset based on a novel inductive generation pipeline. Training is performed using adapters Houlsby et al. (2019) to improve unstructured evidence citation and mitigate the lost in the middle problem. For the latter, previous work has shown that fine-tuning with data augmentation (e.g., shuffling documents; Zhang et al., 2024b) can help achieve this. Given this, we construct SUnsET so that documents are modular: documents are broken down into discrete sections, so that data augmentation through shuffling document sections (thus shuffling global structure) is possible. We first present the inductive pipeline approach used to generate SUnsET, followed by our two fine-tuning schemes.
3.1 Generating SUnsET
Our pipeline generates long documents paired with queries, and summaries which address those queries. Each summary additionally includes citations which reference relevant text spans in the original document. We make several design decisions intended to overcome known problems in synthetic data generation, including the potential for low diversity Honovich et al. (2023); Wang et al. (2023) and labeling errors Chen et al. (2024). This includes taking a six stage pipeline approach which generates synthetic data inductively, and validation steps which refine summaries, refine evidence, and reject bad summaries and evidence.
@@ 3.1 Generating SUnsET
1]] Writing the unwritable requires a recognition of the limitations of language, and a willingness to push against those boundaries.
@@
[…]
Cost Comparison
Manually annotating data of the kind in SUnsET is highly expensive, requiring annotators to read long sets of documents with long summaries and verifying the quality of the references. As a comparison, SQuALITY Wang et al. (2022) is a similar dataset to ours in terms of document and response size, and they paid Upwork workers \$13 to write each response, followed by \$8 to review each response in their data. As we generated 11,309 responses in SUnsET, this alone would have cost \$237,468. In contrast, generating SUnsET, including documents, questions, responses, and evidence, cost around \$200.
SUnsET Non-Pipelined Title + Doc
Metric Q S D Q S D Q S D
TTR 0.75 0.84 0.82 0.67 0.80 0.35 0.63 0.78 0.35
Cos 0.81 0.73 0.68 0.73 0.72 0.04 0.66 0.61 0.04
Len 13.45 226.5 3767.4 9.85 23.79 474.8 10.21 24.45 433.8
Table 1: Statistics and diversity metrics of synthetic data. Metrics are average type-token ratio (TTR) Bestgen (2023), embedding cosine distance (Cos), and average word length (Len). Columns differentiate between (Q)uestion, (S)ummary and (D)ocument metrics in each dataset. Bold is highest diversity across datasets.
Evaluation
[…]
Dataset Topic Diversity
Non-Pipelined 0.506
Title + Doc 0.356
SQuALITY (human, stories) 0.705
LexAbSumm (human, legal text) 0.673
ScholarQABench (human, scientific docs) 0.695
SUnsET 0.679
Table 2: Topic diversity scores using the approach from Terragni et al. (2021). Shading indicates magnitude of diversity score.
To validate SUnsET diversity, we generate two baseline datasets. The first is generated by combining all the steps in Figure 3 into one prompt, forcing the model to simultaneously perform all tasks to generate each example (called Non-Pipelined). The second includes a title generation step to seed each document (called Title + Doc, see Figure 18 in Appendix A for prompts). We compare each dataset using samples of 100 documents along lexical and semantic diversity metrics in Table 1. Further, in Table 2 we compare the topic diversity (following Terragni et al. 2021) between these datasets, as well as three human-written datasets: SQuALITY Wang et al. (2022), LexAbSumm Santosh et al. (2024), and ScholarQABench Asai et al. (2024), (see Appendix C. Our approach generates longer documents with longer summaries than baseline non-pipelined approaches, which also tend to be much more diverse. Additionally, our pipeline produces documents with topic diversity similar to that of human written datasets.
@@ 3.2 Training Complementary Adapters
Previous work has demonstrated that altering the position embeddings of LLMs either directly or through fine-tuning can help to overcome positional biases Hsieh et al. (2024); Zhang et al. (2024b). We design SUnsET documents so that they are modular, having global coherence at the level of the full document and local coherence at the level of discrete sections. Given this, we experiment with position-aware and position-agnostic training in order to observe their impact on evidence selection and quality, as well as summary quality. For position-aware training, we concatenate all document sections together in their natural order to construct the context, while for position-agnostic training, we shuffle the document sections before concatenating them, thus randomizing the global structure of the position embeddings while maintaining the local structure. This gives us two adapters for each model in our experiments. The prompt we use for training is provided in Appendix A Figure 19, and all training is performed using supervised fine-tuning on SUnsET data using LoRA Hu et al. (2022). In all cases we fine tune using the Huggingface Transformers implementation of LoRA Hu et al. (2022) with a rank and $\alpha$ of 16 applied to all linear operators of each model.